A little demonstration of ε-TEX

Jörg Knappen Zentrum für Datenverarbeitung Johannes Gutenberg-Unversität Mainz D-55099 Mainz Allemagne

          E-Mail: knappen@vkpmzd.kph.uni-mainz.de

August 28, 1997

Abstract

The most new primitives of ε-TEX are mostly made for package writers and TEX programmers. However, there is one notable exception, the new \middle primitive which extends the well known pair \left and \right. This paper shows some nice formulæ, and how they can be typeset using ε-TEX and the new \middle primitive.

The new \middle primitive may only be used between a \left and a corresponding \right. It provides a delimiter which grows to the same height as the surrounding left and right delimiters and has by default a tight spacing to both sides. This default behaviour allows us to set the following formula from the bottom of p. 147 of the TEXbook

\frac{a + 1}{b}/\frac{c + 1}{d}

as follows:

\begin{displaymath}
\left.\frac{a+1}{b}\middle/\frac{c+1}{d}\right.
\end{displaymath}

Other examples require a looser spacing around the middle delimiter, as the two other examples on the same page suggest.

(x ∈ A(n) | x ∈ B(n))

and

⋃_{n} X_{n} ∥ ⋂_{n} Y_{n}

are typeset

\begin{displaymath}
\left(x\in A(n) \middleop| x\in B(n)\right)
\end{displaymath}
\begin{displaymath}
\textstyle \left.\bigcup_n X_n \middleop\| \bigcap_n Y_n\right.
\end{displaymath}

using the macro \middleop which is defined as

\newcommand*\middleop[1]{\;\middle#1\;}

There can be more than one \middle between a \left and a \right, as the following example demonstrates

⟨Ψ̄ | Ô | Φ⟩

This example is entered as

\begin{displaymath}
\left< \bar{\Psi} \middleop| \hat O \middleop| \Phi \right>
\end{displaymath}

\begin{matrix}c \\ ◯ \end{matrix} Copyright 1996 Jörg Knappen. All rights reserved.